22. A coil has an inlet temperature of 50 oF and outlet of 85 oF. If the mean temperature of coil is 100 oF, find

the bypass factor of the coil. (SHOW THE SOLUTION)


A. 0.4 B. 2 C. 1.667 D. 0.3


1
Expert's answer
2022-04-14T04:27:03-0400

For heating coil t3 < t2 < t1

t1 = 50

t2 = 85

t3 = 100

BPF = (t3 - t2) / (t3 - t1 ) = (100 - 85) / (100 - 50) = 0.3